OpenStudy (lilmike234):

Which of the following equations represents a degenerate ellipse? 9x^2 +4y^2 = 1 x^2 y^2 ---- + ---- = 2 14 19 x^2 y^2 ----- = (-) ---- 10 18 All of the above are degenerate ellipses.

OpenStudy (loser66):

What do you think?

OpenStudy (lilmike234):

The third option

OpenStudy (loser66):

why?

OpenStudy (lilmike234):

I honestly just guessed

OpenStudy (loser66):

ok, the first one is the formula of a circle \((3x)^2+(2y)^2=1\) That shows the last one is NOT the correct answer.

OpenStudy (loser66):

The third one is \(\dfrac{x^2}{10}=-\dfrac{y^2}{18}\) add \(\dfrac{y^2}{18}\) both sides, what do you get?

OpenStudy (lilmike234):

y^2 ?

OpenStudy (loser66):

nope it becomes \(\dfrac{x^2}{10}+\dfrac{y^2}{18}=0\) The standard formula of an ellipse has the right hand side =1, NOT =0, hence the third option is NOT the correct answer.

OpenStudy (lilmike234):

So it would be the first option right ?

OpenStudy (loser66):

the second one.

OpenStudy (lilmike234):

I was right it was the third option